OpenStudy (anonymous):

6x-2(4-3x)=7-3(17-x)

OpenStudy (anonymous):

Anyone reply please..:-(

OpenStudy (anonymous):

Wanna work on this together again?

OpenStudy (anonymous):

Ceratinly...

OpenStudy (anonymous):

Awesome. What do we do first?

OpenStudy (anonymous):

I just wanna clarify,

OpenStudy (anonymous):

I ve worked it out...

OpenStudy (anonymous):

Here it goes...

OpenStudy (anonymous):

Okie dokie.

OpenStudy (anonymous):

6x-(8-6x)=7-51-3x

OpenStudy (anonymous):

Close

OpenStudy (anonymous):

6x-8-6x=7-51-3x 6x-6x-8=-44-3x

OpenStudy (anonymous):

6x-6x+3x=-44+8 3x=-36

OpenStudy (anonymous):

x=-36/3 x=-12

OpenStudy (anonymous):

6x-2(4-3x)=7-3(17-x) 6x-(8-6x)=7-51+3

OpenStudy (anonymous):

is that perfect..?

OpenStudy (anonymous):

On your first simplification, do you see where it went wrong?

OpenStudy (anonymous):

Well, look back at your first simplification.

OpenStudy (anonymous):

No can u please guide me...

OpenStudy (anonymous):

Sure thing!

OpenStudy (anonymous):

The original problem was: 6x-2(4-3x)=7-3(17-x) You simplified to 6x-(8-6x)=7-51-3x

OpenStudy (anonymous):

Everything was good except one part. What is -3 times -x?

OpenStudy (anonymous):

It's 3x

OpenStudy (anonymous):

Yup. So what do we need to change?

OpenStudy (anonymous):

Reminder: The original problem was: 6x-2(4-3x)=7-3(17-x) You simplified to 6x-(8-6x)=7-51-3x

OpenStudy (anonymous):

6x-(8-6x)=7-51+3x

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

And now simplify the parentheses on the left.

OpenStudy (anonymous):

6x-8+6x=7-51+3x

OpenStudy (anonymous):

perfecto!

OpenStudy (anonymous):

Now combine similar terms.

OpenStudy (anonymous):

6x+6x-3x=7-51+8

OpenStudy (anonymous):

good now squish them together.

OpenStudy (anonymous):

9x=-36 x=-36/9 x=-12

OpenStudy (anonymous):

You were right up to here: x=-36/9

OpenStudy (anonymous):

36/ three would be 12.

OpenStudy (anonymous):

So just recheck the -36/9 part and you'll be all set!

OpenStudy (anonymous):

Sorry simple mistake, it;'s -4

OpenStudy (anonymous):

Perfect! That's the correct answer when you plug it back in!

OpenStudy (anonymous):

Sure, thanks, r u a teacher...

OpenStudy (anonymous):

Nope, just a calculus student. ;) You asked me earlier. lol

OpenStudy (anonymous):

Excellent:-)

OpenStudy (anonymous):

Bye for now enjoy...